OpenStudy (anonymous):

Rationalize the numerator. (sqrt19-(7))/(sqrt14+(7))

jhonyy9 (jhonyy9):

do you know what mean ,,rationalize" ?

jhonyy9 (jhonyy9):

so in this case the numerator ... this is sure ? not the denominator ?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

It is the numerator

jhonyy9 (jhonyy9):

ok so than you need rationalizing the numerator do you know what mean ,,rationalizing " ?

OpenStudy (anonymous):

No

jhonyy9 (jhonyy9):

so do you know that a^2 -b^2 = (a-b)(a+b) yes ?

OpenStudy (anonymous):

And it is only sqrt19 and sqrt14 that is why I tried to separate them out. Hope I did it right.

OpenStudy (anonymous):

yes

jhonyy9 (jhonyy9):

so yes you have do it right than there are sqrt19 -7 ---------- = sqrt14 +7 so because you know that a^2 -b^2 =(a-b)(a+b) and you need rationalizing the numerator what is in this case sqrt19 -7 so than you need to know that sqrt19 -7 has the conjugate sqrt19 +7 yes ?

OpenStudy (anonymous):

yes

jhonyy9 (jhonyy9):

hopw that you see it now that if you multiplie the numerator by conjugate you will get for numerator (sqrt19 -7)*sqrt(19 +7) = 19 -49 = -30 but this is very important that you can multiplie the numerator by the conjugate just you will multiplie with this conjugate of numerator the denominator too so than you will write (sqrt19 -7) *(sqrt19 +7) 19 -49 ---------------------- = ------------ (sqrt14 +7) *(sqrt19 +7) can you ending the calcule ?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

I am an idiot when it comes to math

jhonyy9 (jhonyy9):

ok so than for numerator there will be -30 ,yes ?

OpenStudy (anonymous):

yes

jhonyy9 (jhonyy9):

ok and for denominator there are (sqrt14 +7) *(sqrt19 +7) = sqrt(14*19 +7sqrt14 +7sqrt19 +49 = sqrt266 + 7(sqrt14 +sqrt19) +49 so can you simplifie the sqrt266 ?

OpenStudy (anonymous):

would the answer be -30/sqrt266+7sqrt19+7sqrt14+49

jhonyy9 (jhonyy9):

yes but you see that from7sqrt19 and 7sqrt14 you can factorizing out the 7

OpenStudy (anonymous):

or -30/sqrt266+14sqrt19+49

jhonyy9 (jhonyy9):

how you have got 14sqrt19 ? and where will be the sqrt14 ?

OpenStudy (anonymous):

that's why I think it is -30/sqrt266+7sqrt19+7sqrt14+49

jhonyy9 (jhonyy9):

so than this will be the right answer sure ok ? than good luck bye was my pleasure to help you

OpenStudy (anonymous):

thank you. What did you mean about factoring out 7

jhonyy9 (jhonyy9):

so than you will get 7sqrt19 +7sqrt14 = 7(sqrt19 +sqrt14)

OpenStudy (anonymous):

so i guess either would be ok

jhonyy9 (jhonyy9):

yes good luck bye